OpenStudy (anonymous):

Solve by Factoring. x^3-5x^2-4x+20=0

OpenStudy (anonymous):

\[x= [5,2,-2]\]

jhonyy9 (jhonyy9):

x2(x-5)-4(x-5)=(x2 -4)(x-5)=0 x2 -4=0 x2 =4 x_1,_2= +/- 2 x-5=0 x_3=5 - so those solutions are -2,2,5